Let p be an odd prime number, so that p = 2k + 1 for some positive integer k. Prove that
(k!)2 ≡ (−1)k+1 mod p.
Hint: Try to see how to group the terms in the product
(p − 1)! = (2k)! = 1 ∗ 2 ∗ 3 · · ·(2k − 2) ∗ (2k − 1) ∗ (2k)
to get two products, each equal to k! modulo p.
1
Expert's answer
2019-10-21T12:34:39-0400

Consider the product "(p-1)! = 1\cdot(p-1)\cdot 2\cdot (p-2)\cdots k\cdot (p-k)"



"1\cdot(p-1)\cdot 2\cdot (p-2)\cdots k\cdot (p-k)\ \equiv\ 1\cdot (-1)\cdot 2\cdot (-2)\cdots k\cdot (-k)\ \equiv\ -1 \pmod{p}"

(See Wilson's theorem).


"1\cdot (-1)\cdot 2\cdot (-2)\cdots k\cdot (-k) = (-1)^k (k!)^2"

Therefore:


"(-1)^k (k!)^2 \ \equiv\ -1 \pmod{p}""(k!)^2 \ \equiv\ (-1)^{k+1} \pmod{p}"